Mr brown is 3 times as old as Sam. The difference in their ages is 22 how old is each
Found 2 solutions by fractalier, Alan3354:
Answer by fractalier(6550)

Call Sam's age, x.
Then Mr. Brown's age must be 3x.
Since their difference is 22, we can write
3x - x = 22
2x = 22
x = 11 (Sam's age)
3x = 33 (Mr. Brown's age)

Answer by Alan3354(69443)

Mr brown is 3 times as old as Sam. The difference in their ages is 22 how old is each
--------------
3x as old --> the difference is 2x Sam's age.
2S = 22
S = 11
B = 33